OpenStudy (anonymous):

**WILL FAN AND MEDAL** If f(x) = x2 and g(x) = x + 6, find g(f(0)).

OpenStudy (anonymous):

X^2 or 2x?

OpenStudy (anonymous):

my bad x^2

Vocaloid (vocaloid):

first, if f(x) = x^2, then what is f(0)?

OpenStudy (anonymous):

that's the question...

Vocaloid (vocaloid):

what is the value of f(0)?

OpenStudy (anonymous):

0?

Vocaloid (vocaloid):

right, f(0) = 0 g(x) = x + 6 g(f(0)) = g(0) = ?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

nope, please try again g(x) = x + 6 what is g(0)?

OpenStudy (anonymous):

6

Vocaloid (vocaloid):

yup, that's it

OpenStudy (anonymous):

thank you!